### v1

@@ 1 Introduction
Figure 1: A comparison of AAL Task (left) and RAVAR task (right). The AAL task requires either user-guided or automatically generated Regions Of Interest (ROIs), relying on human postprocessing to select the individual of interest after obtaining the predictions for all individuals. In contrast, RAVAR relies on textual references on specific individuals, improving efficiency. RAVAR requires the model to automatically identify the location and predict atomic actions for the individual of interest.
@@
To establish an initial benchmark, we assess $11$ well-established approaches from multiple related research domains, e.g., video-based Atomic Action Localization (AAL), Video Text Retrieval (VTR), and Video Question Answering (VQA). Our experiments reveal that none of the benchmarked methods delivers sufficient performance. For example, previous methods within the domains of VQA and VTR typically yield predictions of a coarse analysis regarding actions, whereas strategies in AAL face challenges in spatially localizing the correct referring individual. We attribute this phenomenon to the massive irrelevant information provided by the visual data, distracting the model from the referring location. How to suppress the referring irrelevant information becomes a critical challenge to pursue better RAVAR performances.
To address this, we introduce the new RefAtomNet approach specialized in simultaneously solving the tasks of understanding textual referring expressions and using them as guides for spatial localization and classification of atomic actions. We begin by computing textual, visual, and newly proposed location-semantic tokens using a large foundation model and well-established object detector. The latter stream is designed to harvest semantic cues of different scene entities: it leverages a pre-trained object detector to localize and identify objects, after which the location-semantic tokens are computed as a concatenation of visual bounding box embeddings and text embeddings of the object category. Another novel aspect of our work is the introduction of cross-stream agent attention and agent token fusions, which builds on the concept of agent attention [25] and re-defines it for cross-stream compatibility: agent tokens are reformulated into a 1D sequence format through the use of fully connected layers, bypassing the initial need for 2D pooling, removing the depthwise convolution branch, and forgoing 2D positional encodings. This mechanism enhances the model’s ability to filter and emphasize relevant information from multiple data sources effectively. RefAtomNet achieves the highest performance for RAVAR, while a detailed ablation study demonstrates the effectiveness of each component.
@@ 2 Related Work
Referring Scene Understanding. Referring scene understanding aims to locate parts of interest within images or videos guided by natural language, exhibiting the utility in many computer vision tasks including autonomous driving [81] and video editing [6]. The development of this field cannot be advanced without the contribution of high-quality open-source datasets and benchmarks [2, 87, 75, 28, 68, 81, 12, 45]. For example, the CLEVR-Ref+ benchmark is proposed by Liu et al. [50] to achieve visual reasoning with referring expressions. Li et al. [43] proposed referring image segmentation by using a recurrent refinement network. Wu et al. [81] proposed referring multi-object tracking benchmark. However, there are no referring understanding works focusing on atomic video action analysis, whereupon we first conduct RAVAR benchmarking with our RefAVA dataset.
Video Text Retrieval. Video Text Retrieval (VTR) aims at matching relevant video content with text. Amidst the rise of Vision-Language foundation models like CLIP [63] and BLIP [40], efforts [77, 52, 53, 39, 91, 82, 9, 54, 47, 73] are being made to apply powerful pre-trained models for promoting their competencies in VTR. XCLIP [53] introduced multi-grained contrastive learning for end-to-end VTR, by counting all the video-sentence, video-word, sentence-frame, and frame-word contrasts. MeVTR [91] achieved promising performance in VTR via integrating key event video representation alongside a novel MeVTR loss. BLIPv2 [39] pre-trained a Querying Transformer (QFormer) bootstrapping vision-and-language representation learning and vision-to-language generative learning. The primary distinction between VTR and RAVAR lies in their inference phase tasks: VTR involves querying using either text or video to localize respective elements, whereas RAVAR necessitates integrating both textual and visual data to achieve fine-grained subject retrieval and atomic action prediction in videos.
@@
Atomic Video Action Recognition and Localization. Atomic video-based action recognition [22, 11] and localization [23] involves the identification and analysis of the most fundamental, indivisible actions or movements performed by humans for single and multiple-person scenarios. Compared with the general video action recognition task, atomic video-based action localization is much more fine-grained and is always formulated in a multi-label manner with bounding box predictions. Most of the existing convolutional neural networks (CNN) [4, 76, 16, 17] and the transformer [65, 76, 44, 78, 80, 59, 22] networks for human action recognition are commonly used in the atomic video-based action localization by changing the classification head into multi-label manner, adding additional bounding box prediction head, and integrating region of interest features from human detector. Ryali et al. [65] proposed a hierarchical vision transformer with high efficiency and precision within the realm of the existing methods using video as input. Wang et al. [76] proposed scaling video-masked autoencoders with dual masking.
Current methods for predicting individual actions in multi-person scenarios either require manual input to isolate individuals through video cropping for atomic action recognition or generate predictions of all detected individuals for atomic action localization [61, 79, 64], necessitating further human selection effort and impairing practicality. Most of the existing AAL methods are not particularly designed for the RAVAR task. We thereby introduce RefAtomNet, a novel method that utilizes location semantics atop predicted location and scene semantic information derived from the scenario together with the textual reference and visual cues, then further uses cross-stream agent attention and agent token fusions to suppress redundant information.
@@ 3 RAVAR: Established Benchmark
Table 1: An overview of the referring scene understanding datasets and our RefAVA dataset; Our task is Referring Atomic Video Action Recognition (RAVAR), whereas existing benchmarks focus on Referring Object Detection (ROD), Refering Video-based Object Segmentation (RVOS), and Referring Multi-Object Tracking (RMOT).
@@ 3.1 Introduction of the RefAVA Dataset
Textual Annotations. To acquire precise textual annotations for the individuals of interest, $7$ annotators manually provided the textual annotations according to the key frame bounding boxes presented in the AVA dataset [23]. Cross-checking among all the annotators for the annotated individuals is conducted to deliver high textual annotation quality, which ensures an accurate description of the instance during the annotation process to the utmost extent. The individuals with high disagreements from the annotators are removed from the dataset.
Dataset. We selected $17,946$ video clips from $127$ movies of the AVA dataset, preserving the most complex scenarios, and annotated each person on each center frame of every $90$s video clip. In total, RefAVA has $36,630$ labeled instances, and we split them into $22,658$ train instances, $10,916$ validation instances, and $3,056$ test instances, where the samples from different sets are from different movie scenarios. The textual annotations cover information on approximate age, gender, appearance, relative position in the center frame, etc., and without action description. The chord visualization in Fig. 2 reveals the quantitative causal relations between the most frequent notional words in different categories from the textual references. Further, Fig. 2 shows the statistics of instance amount towards the annotated person number inside the scenario where each instance belongs to. We deliver the comparison of the statistics between representative existing referring scene understanding datasets and our RefAVA dataset in Tab. 1. The atomic actions involve $80$ categories covering Object Manipulation (OM), Person Interactions (PI), and Person Movement (PM).
@@ 3.2 The Baselines for the RAVAR Benchmark
3.3 Metrics.
We use multi-label mean Average Precision (mAP), Area Under the Receiver Operating Characteristic curve (AUROC), and the mean Intersection over Union (mIOU) as metrics, of which mIOU has less priority since we prioritize the atomic video action recognition task. The conventional atomic action localization task evaluates the localization and the atomic action predictions jointly by choosing a threshold of $mIOU{=}0.5$ to calculate mAP, however, we evaluate these two terms separately for a better illustration of the particular capability in RAVAR. More details are delivered in the supplementary.
@@ 4.1 Overview
We propose a new model, RefAtomNet, with an overview provided in Fig. 3. It leverages three token streams: visual, textual reference, and location-semantic streams. Visual tokens are extracted from the video using ViT [15], while textual reference tokens are obtained from the text using BERT [14]. Both streams are enhanced by QFormer [39]. Additionally, in the location-semantic stream, location-semantic tokens are extracted from the center frame by fusing predicted object coordinates and semantic embeddings of the object categories estimated by a frozen object detector and BERT [14]. These tokens contribute by incorporating location-semantic embeddings of the individual scene entries. To suppress irrelevant information for each stream, we introduce a 1D sequential agent attention inspired by Han et al. [25]. Agent attention relies on agent tokens, i.e., tokens extracted by a pooling layer. Unlike their 2D approach for images, we extract agent tokens using fully connected layers on the input tokens, similar to the extraction of Query, Key, and Value in transformers. We further propose the agent-based location-semantic aware attentional fusion mechanism to achieve better amplification of relevant information during the cross-stream exchange. The designed agent-based location-aware attentional fusion mechanism guides the model’s attention toward the most relevant visual tokens, leading to improved visual-textual alignment. The detailed structure is discussed in the following subsections.
@@ 4.2 RefAtomNet
Figure 3: RefAtomNet operates on three streams: visual, textual reference, and location-semantic aware streams. It first employs agent attention for self-enhancement of the latter two streams. Then, RefAtomNet uses cross-stream agent token fusion and cross-stream agent attention fusion to suppress irrelevant tokens in the visual stream.
Establishing a meaningful connection between video content and textual descriptions is vital in our task. We, therefore, introduce cross-stream agent attention and token fusion – an approach for improving information exchange of the three streams, i.e., textual reference, visual, and location-semantic streams.
Agent Tokens and Sequential Processing. An important element in our approach is the use of agent tokens, which is inspired by the recently proposed agent attention [25]. Derived through fully connected layers from input tokens—akin to the extraction of Query, Key, and Value in transformers—agent tokens aggregate essential information via agent-key and agent-query pairs. To enable the use of agent attention for 1D sequential tokens resulting from our streams, we redefine the 2D agent spatial tokens, proposed in [25], into a 1D sequential format to enable multi-stream fusion. Specifically, we replace the 2D pooling in agent token generation with linear projection. We also eliminate the depthwise convolutional branch and biased position encoding to better suit agent acquisition in our model. These 1D sequential agent tokens serve a dual function. First, they achieve the aggregation of the crucial information from corresponding agent-key and agent-query pairs. Secondly, they efficiently redistribute the aggregated information back to the original value tokens, enhancing the attention mechanism’s focus and reducing irrelevant cues.
@@
where $FFN$ indicates Feed Forward Network. These steps ensure precise model attention, enhancing the contextual relevance of the resulting tokens.
@@
Then, we calculate the cross-stream irrelevance-suppressed attention for the agent query attention of visual stream through Eq. 8 to achieve cross-stream agent attention fusion to suppress irrelevant information in the agent query attentions of the visual stream, the operations used in the second and third terms are abbreviated as C-ATT and T-ATT in Fig. 3,
@@ 5.1 Implementation Details
We conduct experiments on four NVIDIA A100 GPUs. We use BertAdam [30] optimizer with learning rate $lr=1e^{-4}$, batch size $128$, learning rate decay $0.9$, and warmup ratio $0.1$. The agent number and head number in our RefAtomNet are $4$ and $1$. Our model has $214M$ trainable parameters. The weight for $L_{MSE}$ and $\alpha$ are $5$ and $0.125$.
@@ 5.2 Experimental Results
Main Results. Results of the main experiments on the proposed RefAVA benchmark are summarized in Tab. 2. We compare our RefAtomNet model with a multitude of published approaches adopted to suit our task (see Section  3.2) stemming from the fields of (1) Atomic Action Localization (AAL); (2) Video Question Answering (VQA); and (3) Video-Text Retrieval (VTR). Baseline methods from the AAL group consistently underperform in spatially localizing the action according to the textual reference (measured as mIOU). Even though the ability of the human spatial localization is not the major expected output, it can still generally illustrate if the model can grasp the correct referring person or not. Among the AAL methods, X3D [16] achieves the highest AUROC score of $59.09\%$ and $64.51\%$ on the val and test sets, respectively. The VQA and VTR baselines work better than the AAL baselines since they can reasonably capture the referring person while delivering acceptable atomic human action recognition performances, which benefits from the text-aware pretraining of these two tasks. AskAnything [41] from the VQA group achieves $20.09\%$ and $22.35\%$ of mIOU, $51.42\%$ and $52.25\%$ of mAP, and $66.12\%$ and $69.35\%$ of AUROC on the val and test sets. BLIPv2 [40] from the VTR group delivers $32.99\%$ and $32.75\%$ of mIOU, $52.13\%$ and $53.19\%$ of mAP, and $66.56\%$ and $69.92\%$ of AUROC, on the val and test sets, respectively. Our RefAtomNet achieves state-of-the-art performances, outperforming the best baseline BLIPv2 [39] by $5.23\%$, $3.85\%$, $3.17\%$ and $3.67\%$, $4.33\%$, $4.03\%$ of mIOU, mAP, and AUROC, on val and test sets.
Table 2: Experimental results on our RAVAR benchmark. We leverage mIOU to evaluate the localization for the person of interest, and mAP and AUROC to evaluate the referring atomic action recognition, where mAP and AUROC are prioritized over mIOU. Baselines are selected from three relevant domains, i.e., AAL, VQA, and VTR.
@@
Ablations of the Individual Modules. In Tab. 3, we show the ablation for the components of RefAtomNet by removing each of the proposed designs, where LSAS indicates the location-semantic stream, CAAF indicates the cross-stream agent attention fusion, CATF indicates the cross-stream agent token fusion, and w/o ALSAF indicates by simply using addition to fuse these three streams and without the agent-based location-semantic aware attentional fusion mechanism. Compared with the ablation w/o ALSAF, RefAtomNet achieves promising improvements of $10.92\%$, $5.28\%$, $4.42\%$ and $7.33\%$, $6.26\%$, $5.75\%$ in terms of mIOU, mAP, and AUROC for the val and test sets, respectively. It indicates that using simple aggregation of the three streams makes the model distract on the important cues for the referring person, illustrated by the large decay of mIOU metric. Compared with the ablation w/o LSAS, we find that the location-semantic tokens benefit more for the localization of the correct person in the center frame. The advantages of the LSAS on atomic video action recognition metrics are highlighted in the test set which has higher scenario diversity compared with the val set. Using CAAF and CATF, each of them brings promising benefits by suppressing the irrelevant information in the visual tokens.
Comparison with Other Fusion Mechanisms. We compare our proposed agent-based location-semantic aware attentional fusion mechanism with late fusion addition, multiplication, concatenation, AttentionBottleNeck [55], and McOmet [94] in Tab. 3. Most of the existing fusion strategies cannot deal with the RAVAR task well due to the lack of concentration on informative cues. For addition, multiplication, McOmet [94], and concatenation, the mIOU is less than $30\%$, since these approaches do not prioritize the most relevant visual tokens, which will lead to an unexpected matching between the reference and the visual cues. AttentionBottleNeck [55] alleviates this issue by highlighting different tokens individually through the self-attention mechanism. Our approach outperforms the AttentionBottleNeck [55] by large margin through efficiently suppressing those irrelevant visual tokens on the agent tokens and agent attention masks, resulting in $4.75\%$, $5.01\%$, $4.66\%$, and $3.40\%$, $3.50\%$, $2.87\%$ improvements in terms of mIOU, mAP, and AUROC on the val and test sets, respectively.

### v2

@@ 1 Introduction
Figure 1: A comparison of AAL Task (left) and RAVAR task (right).
@@
To establish an initial benchmark, we assess $15$ well-established approaches from multiple related research domains. Our experiments reveal that none of the benchmarked methods delivers sufficient performance. For example, previous methods within the domains of VQA and VTR typically yield predictions of a coarse analysis regarding actions, whereas strategies in AAL face challenges in spatially localizing the correct referring individual. We attribute this phenomenon to the massive irrelevant information provided by the visual data, distracting the model from the referring location. How to suppress the referring irrelevant information becomes a critical challenge to pursue better RAVAR performances.
To address this, we introduce the new RefAtomNet approach specialized in simultaneously solving the tasks of understanding textual referring expressions and using them as guides for spatial localization and classification of atomic actions. We begin by computing textual, visual, and newly proposed location-semantic tokens using a large foundation model and well-established object detector. The latter stream is designed to harvest semantic cues of different scene entities: it leverages a pre-trained object detector to localize and identify objects, after which the location-semantic tokens are computed as a fusion of visual bounding box coordinates and text embeddings of the object category. Another novel aspect of our work is the introduction of cross-stream agent attention and agent token fusions, which builds on the concept of agent attention [26] and re-defines it for cross-stream compatibility: agent tokens are reformulated into a 1D sequence format through the use of fully connected layers, bypassing the initial need for 2D pooling, removing the depthwise convolution branch, and forgoing 2D positional encodings. This mechanism enhances the model’s ability to filter and emphasize relevant information from multiple data sources effectively. RefAtomNet achieves the highest performance for RAVAR, while a detailed ablation study demonstrates the effectiveness of each component.
@@ 2 Related Work
Referring Scene Understanding. Referring scene understanding aims to locate parts of interest within images or videos guided by natural language, exhibiting the utility in many computer vision tasks including autonomous driving [87] and video editing [6]. The development of this field cannot be advanced without the contribution of high-quality open-source datasets and benchmarks [2, 93, 81, 30, 73, 87, 12, 48]. For example, the CLEVR-Ref+ benchmark is proposed by Liu et al. [53] to achieve visual reasoning with referring expressions. Li et al. [46] proposed referring image segmentation by using a recurrent refinement network. Wu et al. [87] proposed referring multi-object tracking benchmark. However, there are no referring understanding works focusing on atomic video action analysis, whereupon we first conduct RAVAR benchmarking with our RefAVA dataset. We further clarify that Referring Video Object Segmentation (RVOS) task [54, 80, 21, 59] differs us from multiple perspectives, e.g., RVOS includes action name in input textual reference.
Video Text Retrieval. Video Text Retrieval (VTR) aims at matching relevant video content with text. Amidst the rise of Vision-Language foundation models like CLIP [68] and BLIP [43], efforts [83, 56, 57, 42, 97, 88, 9, 58, 50, 78] are being made to apply powerful pre-trained models for promoting their competencies in VTR. XCLIP [57] introduced multi-grained contrastive learning for end-to-end VTR, by counting all the video-sentence, video-word, sentence-frame, and frame-word contrasts. BLIPv2 [42] pre-trained a Querying Transformer (QFormer) to bootstrap vision-and-language representation learning and vision-to-language generative learning. The key difference between VTR and RAVAR lies in their inference tasks: VTR queries text or video to localize elements, while RAVAR integrates textual and visual data for fine-grained subject retrieval and atomic action prediction in videos.
@@
Atomic Video Action Recognition and Localization. Atomic video-based action recognition [23, 11] and localization [24] involves the identification and analysis of the most fundamental, indivisible actions or movements performed by humans for single and multiple-person scenarios. Compared with the general video action recognition task, atomic video-based action localization is much more fine-grained and is always formulated in a multi-label manner with bounding box predictions. Most of the existing convolutional neural networks (CNN) [4, 82, 16, 17] and the transformer [70, 82, 47, 84, 86, 64, 23] networks for human action recognition are commonly used in the atomic video-based action localization by changing the classification head into multi-label manner, adding additional bounding box prediction head, and integrating region of interest features from human detector. Ryali et al. [70] proposed a hierarchical vision transformer with high efficiency and precision within the realm of the existing methods using video as input. Wang et al. [82] proposed scaling video-masked autoencoders with dual masking. Current methods for predicting individual actions in multi-person scenarios often require manual video cropping for atomic action recognition or generate predictions for all detected individuals, necessitating further human selection and reducing practicality [66, 85, 69]. Most existing AAL methods are not specifically designed for the RAVAR task. We thereby introduce RefAtomNet, a novel method that utilizes location semantics atop predicted location and scene semantic information derived from the scenario together with the textual reference and visual cues, then further uses cross-stream agent attention and agent token fusions to suppress redundant information.
@@ 3 RAVAR: Established Benchmark
Table 1: An overview of the referring scene understanding datasets and our RefAVA dataset; Our task is Referring Atomic Video Action Recognition (RAVAR), whereas existing benchmarks focus on Referring Object Detection (ROD), Referring Video-based Object Segmentation (RVOS), and Referring Multi-Object Tracking (RMOT).
@@ 3.1 Introduction of the RefAVA Dataset
Textual Annotations. To acquire precise textual annotations for the individuals of interest, $7$ annotators manually provided the textual annotations according to the key frame bounding boxes presented in the AVA dataset [24]. Cross-checking among all the annotators for the annotated individuals is conducted to deliver high textual annotation quality.
Dataset. We selected $17,946$ video clips from $127$ movies of the AVA dataset, preserving the most complex scenarios, and annotated each person on each center frame of every $90$s video clip. In total, RefAVA has $36,630$ labeled instances, and we split them into $22,658$ train instances, $10,916$ validation instances, and $3,056$ test instances, where the samples from different sets are from different movie scenarios. The textual annotations cover information on approximate age, gender, appearance, relative position in the center frame, etc., and without action description. The chord visualization in Fig. 2 reveals the quantitative causal relations between the most frequent notional words in different categories from the textual references. Further, Fig. 2 shows the statistics of instance amount towards the annotated person number inside the scenario where each instance belongs. We deliver the comparison of the statistics between representative existing referring scene understanding datasets and our RefAVA dataset in Tab. 1. The atomic actions involve $80$ categories covering Object Manipulation (OM), Person Interactions (PI), and Person Movement (PM). The videos cover diverse scenarios. The test set is sourced from $26$ movies different from those in the train set ($67$ different movies) and val set ($34$ different movies) to achieve the evaluation of generalizability.
@@ 3.2 The Baselines for the RAVAR Benchmark
SF Baselines. Some Single Frame (SF) baselines are adopted, where SAM [33], DETR [3], and REFCLIP [29] are utilized together with the CLIP [68] feature extraction backbone.
VOS Baseline. We adopt one Video Object Segmentation (VOS) baseline using the encoder of the approach from Su et al. [80] and our prediction head.
@@ 4.1 Overview
We propose a new model, RefAtomNet, with an overview provided in Fig. 3. It leverages three token streams: visual, textual reference, and location-semantic streams. Visual tokens are extracted from the video using ViT [15], while textual reference tokens are obtained from the text using BERT [14]. Both streams are enhanced by QFormer [42]. Additionally, in the location-semantic stream, location-semantic tokens are extracted from the center frame by fusing predicted object coordinates and semantic embeddings of the object categories estimated by a frozen object detector and BERT [14]. To suppress irrelevant information for each stream, we propose the agent-based location-semantic aware attentional fusion to achieve better amplification of relevant information during the cross-stream exchange. The detailed structure is discussed in the following subsections.
@@ 4.2 RefAtomNet
Background of QFormer. We rely on BLIPv2 to extract the visual features, where QFormer is the most essential component. Multiple learnable queries are initialized in QFormer as trainable parameters and interact with input data via Transformer attention mechanisms. Each query selectively attends to input parts, capturing task-specific features and updating based on attention outputs. During training, these queries are optimized for specialized information extraction. Queries then directly contribute to generating outputs.
@@
Figure 3: An overview of the RefAtomNet architecture.
Agent Tokens and Sequential Processing. We use agent tokens, inspired by agent attention [26]. Derived through fully connected layers from input tokens, similar to Query, Key, and Value extraction in transformers, agent tokens aggregate essential information via agent-key and agent-query pairs. To enable agent attention for 1D sequential tokens from our streams, we redefine the 2D agent spatial tokens proposed in [26] into a 1D sequential format for multi-stream fusion. Specifically, we replace 2D pooling with linear projection and eliminate the depthwise convolutional branch and biased position encoding to better suit agent acquisition in our model. These 1D sequential agent tokens serve dual functions: aggregating crucial information from agent-key and agent-query pairs and efficiently redistributing this information to the original value tokens, therefore, enhancing attention focus and reducing irrelevant cues.
@@
where $FFN$ indicates the Feed Forward Network. These steps ensure precise model attention, enhancing the contextual relevance of the resulting tokens.
@@
Then, we calculate the cross-stream irrelevance-suppressed attention for the agent query attention of the visual stream through Eq. 8 to achieve cross-stream agent attention fusion to suppress irrelevant information in the agent query attentions of the visual stream, the operations used in the second and third terms are abbreviated as C-ATT and T-ATT in Fig. 3,
@@ 5.1 Implementation Details
We conduct experiments on four NVIDIA A100 GPUs. We use BertAdam [32] optimizer with learning rate $lr=1e^{-4}$, batch size $128$, learning rate decay $0.9$, and warmup ratio $0.1$. The agent number and head number in our RefAtomNet are $4$ and $1$. Our model has $214M$ trainable parameters. The weight for $L_{MSE}$ and $\alpha$ are $5$ and $0.125$. We trained our model for $40$ epochs on our dataset. The text encoder is frozen during training. We use multi-label mean Average Precision (mAP), Area Under the Receiver Operating Characteristic curve (AUROC), and the mean Intersection over Union (mIOU) as metrics, of which mIOU has less priority since we prioritize the atomic video action recognition task.
@@ 5.2 Experimental Results
Main Results. Results of the main experiments on the proposed RefAVA benchmark are summarized in Tab. 2. We compare our RefAtomNet model with a multitude of published approaches adopted to suit our task (see Section  3.2) stemming from the fields of (1) Atomic Action Localization (AAL); (2) Video Question Answering (VQA); (3) Video-Text Retrieval (VTR); (4) Single Frame (SF) baselines; and (5) Video Object Segmentation (VOS). Baseline methods from the AAL group consistently underperform in spatially localizing the action according to the textual reference (measured as mIOU). Even though the ability of the human spatial localization is not the major expected output, it can still generally illustrate if the model can grasp the correct referring person or not. Among the AAL methods, X3D [16] achieves the highest AUROC score of $59.09\%$ and $64.51\%$ on the val and test sets, respectively. The VQA and VTR baselines work better than the AAL baselines since they can reasonably capture the referring person while delivering acceptable atomic human action recognition performances, which benefits from the text-aware pretraining of these two tasks. AskAnything [44] from the VQA group achieves $20.09\%$ and $22.35\%$ of mIOU, $51.42\%$ and $52.25\%$ of mAP, and $66.12\%$ and $69.35\%$ of AUROC on the val and test sets. BLIPv2 [43] from the VTR group delivers $32.99\%$ and $32.75\%$ of mIOU, $52.13\%$ and $53.19\%$ of mAP, and $66.56\%$ and $69.92\%$ of AUROC, on the val and test sets, respectively. Our RefAtomNet achieves state-of-the-art performances, outperforming the best baseline BLIPv2 [42] by $5.23\%$, $3.85\%$, $3.17\%$ and $3.67\%$, $4.33\%$, $4.03\%$ of mIOU, mAP, and AUROC, on val and test sets.
Table 2: Experimental results on our RAVAR benchmark.
@@
Ablations of the Individual Modules. In Tab. 3, we show the ablation for the components of RefAtomNet by removing each of the proposed designs, where LSAS indicates the location-semantic stream, CAAF indicates the cross-stream agent attention fusion, CATF indicates the cross-stream agent token fusion, and w/o ALSAF indicates by simply using addition to fuse these three streams and without the agent-based location-semantic aware attentional fusion mechanism. Compared with the ablation w/o ALSAF, RefAtomNet achieves promising improvements of $10.92\%$, $5.28\%$, $4.42\%$ and $7.33\%$, $6.26\%$, $5.75\%$ in terms of mIOU, mAP, and AUROC for the val and test sets, respectively. It indicates that using simple aggregation of the three streams makes the model distract from the important cues for the referring person, illustrated by the large decay of mIOU metric. Compared with the ablation w/o LSAS, we find that the location-semantic tokens benefit more for the localization of the correct person in the center frame. The advantages of the LSAS on atomic video action recognition metrics are highlighted in the test set which has higher scenario diversity compared with the val set. Using CAAF and CATF, each of them brings promising benefits by suppressing the irrelevant information in the visual tokens.
Comparison with Other Fusion Mechanisms. We compare our proposed agent-based location-semantic aware attentional fusion mechanism with late fusion addition, multiplication, concatenation, AttentionBottleNeck [60], and McOmet [100] in Tab. 3. Our approach outperforms all by suppressing those irrelevant visual tokens on the agent tokens and attention masks.